Which Among the following (general ) annuities will you choose based on the future value?

Option 1

R= 500

I= 8.00 %

Period= 12

t(years) 2

Compounded: 2

Required: What is the future value ?

Option 2

R= 500

I= 8.00 %

Period= 12

t(years) 2

Compounded: 4

Required: What is the future value ?

Which one is the best option?
Answer #1

FV = R* FVIFA

FVIFA =(1+r) ^n - 1 /r

Option 1

FV = 500*((((1+(8%/2))^(2*2))-1)/(8%/2))=2123.23

Option 2

FV = 500*((((1+(8%/4))^(2*4))-1)/(8%/4))=4291.48

When interest is compounded more number of times it is considered better and thus option 2 will be considered better